OpenStudy (anonymous):

Use a graphing utility to graph the function E(t) = 110 cos(120πt −π/3). What is the period of the function?

OpenStudy (jamesj):

The period of the function is the smallest T > 0 such that E(t+T) = E(t) Follow the procedure I used in the last question I answered for you to find T.

OpenStudy (anonymous):

The period of a*Cos(b*t+c) is 2*Pi/b So 2*Pi/(120*Pi) = 1/60 is the period